OpenStudy (anonymous):

Algebra-Inequalities Question: You have 2 circles, one with the radius r and the other one with Radius R. You wish for the difference in the areas of these two circles to be less than or equal to 5 pi. If r+R=10, what is the maximum difference in the lengths of the radii?

OpenStudy (anonymous):

(R^2)*pi -( r^2)*pi = 5*pi so R = 10 - r (10-r)^2 - r^2= 5 pi reduces to 1 since all have pi

OpenStudy (anonymous):

100-20r + r^2 - r^2 = 5 -20r = -95 r = 19/4

OpenStudy (anonymous):

Oh, okay, I got it now.

OpenStudy (anonymous):

:D